### 28 Dec 2023 v1

# Challenge LLMs to Reason About Reasoning: A Benchmark to Unveil Cognitive Depth in LLMs

Abstract
In this work, we introduce a novel evaluation paradigm for Large Language Models, one that challenges them to engage in meta-reasoning. This approach addresses critical shortcomings in existing math problem-solving benchmarks, traditionally used to evaluate the cognitive capabilities of agents. Our paradigm shifts the focus from result-oriented assessments, which often overlook the reasoning process, to a more holistic evaluation that effectively differentiates the cognitive capabilities among models. For example, in our benchmark, GPT-4 demonstrates a performance ten times more accurate than GPT3-5. The significance of this new paradigm lies in its ability to reveal potential cognitive deficiencies in LLMs that current benchmarks, such as GSM8K, fail to uncover due to their saturation and lack of effective differentiation among varying reasoning abilities. Our comprehensive analysis includes several state-of-the-art math models from both open-source and closed-source communities, uncovering fundamental deficiencies in their training and evaluation approaches. This paper not only advocates for a paradigm shift in the assessment of LLMs but also contributes to the ongoing discourse on the trajectory towards Artificial General Intelligence (AGI). By promoting the adoption of meta-reasoning evaluation methods similar to ours, we aim to facilitate a more accurate assessment of the true cognitive abilities of LLMs.
Challenge LLMs to Reason About Reasoning: A Benchmark to Unveil Cognitive Depth in LLMs
Zhongshen Zeng Chinese University of Hong Kong zszeng23@cse.cuhk.edu.hk                        Pengguang Chen Chinese University of Hong Kong pgchen@cse.cuhk.edu.hk
Haiyun Jiang thanks: Corresponding author. Tencent AI Lab haiyunjiang@tencent.com                        Jiaya Jia thanks: Corresponding author. Chinese University of Hong Kong leojia@cse.cuhk.edu.hk
1 Introduction
Pretrained on trillions of tokens and possessed with billions of parameters, today’s large language model (OpenAI2023GPT4TR; ANTHROPIC2023CLAUDE; touvron2023llama) is capable of generating coherent texts that are grammatically correct and seemingly meaningful. If further enhanced with the instruction tuning and human values alignment (e.g. through reinforcement learning from human feedback ouyang2022training), their outputs are on par with and even more superior than that of human beings. With the hope of differentiating different model’s cognitive ability, math questions are often selected as a proxy evaluation task. However, despite the complexity and diversity of these math problems, recent SOTA LLMs (OpenAI2023GPT4TR; yu2023metamath; Gou2023ToRAAT) have been able to achieve accuracy rates exceeding 80% (luo2023wizardmath) on multi-step math reasoning datasets like GSM8K (cobbe2021training).
Upon a detailed examination of the design principles and objectives of current math datasets, we identified several key shortcomings. Firstly, the majority of these datasets focus on result-oriented metrics, scoring accuracy based solely on the final answer, without considering the underlying reasoning process. With the chain of thought methodology (Wei2022ChainOT) and its derivative techniques (Chen2022ProgramOT; Yao2023TreeOT) emerged as the de facto standard for reasoning processes, we argue that the result-driven evaluation method may be insufficient for a comprehensive assessment of the intended cognitive and reasoning capabilities. Secondly, as a recent study (PasterHungarianTesting) suggests that some LLMs who achieved SOTA performances in GSM8K and MATH (hendrycks2021measuring) benchmarks have unexpectedly low performance when facing newly released Hungarian high school exams. This raises concerns about the data contamination and potential overfitting to the benchmarks. It also challenges the efficacy of these benchmarks in differentiating model capabilities.
Figure 1: Comparative Performance in GSM8K and DiagGSM8K: This graph illustrates the performance of GPT-4, Claude2, and GPT3-5 in the standard GSM8K benchmark versus our novel DiagGSM8K benchmark. While these models show similar performance levels in GSM8K, a marked variance is observed in DiagGSM8K, highlighting the differentiating power of our benchmark in evaluating deeper reasoning capabilities.
Figure 2: Analysis of Training Sets and Reasoning Path Mastery: The left side of the figure demonstrates the typical inclusion of correct reasoning paths in current math reasoning training sets, despite variations focusing on aspects such as reasoning types (e.g. program of thought) or solution diversity. The right side depicts the limitations of models trained exclusively on correct reasoning paths, showing their inability to assess the validity of alternative reasoning paths for the same problem. This highlights a critical gap in the training paradigm, where models demonstrate basic imitation skills but lack a deep understanding of the underlying logical rules, leading to a superficial grasp of reasoning processes.
In response to these identified limitations, we have introduced a novel paradigm, namely challenging LLMs to reason about different reasoning. This approach is specifically designed to augment existing evaluation benchmarks (as illustrated in Figure-3). Under this paradigm, LLMs are tasked to adopt a role akin to that of a teacher, assessing solutions by determining correctness, identifying potential initial errors, and providing reasons for these errors. Following this design principle, we have developed a novel benchmark named DiagGSM8k. This benchmark, characterized by instances manually labeled by experts and rigorously reviewed, serves as a robust tool for a qualitative and quantitative assessment of language models. Our findings indicate that most SOTA models only manage to achieve single-digit accuracy in this more nuanced assessment. As demonstrated in Figure-1, although SOTA models exhibit comparable performance in GSM8K, there is a considerable variance in their effectiveness on our benchmark, with discrepancies up to tenfold.
We argue that our evaluation paradigm does not only provide a more differentiating metric that focuses on the reasoning process over mere computational outcomes, but also sheds light on fundamental deficiencies within the current evaluation and training approaches. As visualized by Figure-2, we observe that recent SOTA models—regardless of whether they are trained on datasets that augment the respective benchmarks in terms of difficulty (luo2023wizardmath), coverage (lee2023platypus; yue2023mammoth), reasoning types (yu2023metamath; yue2023mammoth), solution diversities (yuan2023scaling), solution types (Gou2023ToRAAT; yue2023mammoth), or through reinforcement learning (Uesato2022SolvingMW; luo2023wizardmath), are essentially trained by teaching the language models to output a few solution paths within a large search space. Although these models can generate seemingly correct solutions, their grasp of underlying rationale and principles is often superficial and unsophisticated. This is evidenced by issues such as unit inconsistency during calculations (Toh2023VerityMathAM) which highlights a lack of fundamental ontological understanding, and the inability to discern nuanced differences between various reasoning paths for the same problems it was trained on. These core cognitive understanding abilities, vital for benchmarks to assess, have been overlooked due to the prevalent result-oriented metrics and evaluation paradigms. For downstream applications, particularly in education and consulting, the critical factor for success lies in the capability to rigorously analyze different plausible solutions, offering a comprehensive and well-rounded exploration of the solution space. We will provide a more in-depth discussion in section-5.
The ’reason about reasoning’ principle, central to this benchmark, represents a versatile evaluation approach that can be readily integrated into various existing benchmark suites to enhance their comprehensiveness. Our investigations, detailed in Section-4, uncover significant limitations in the performance of current state-of-the-art models. Furthermore, as discussed in Section-5, we observe that prevalent training paradigms predominantly focus on inductive learning, often failing to cultivate the emergence of underlying rules and principles. This observation underscores the need for a more reflective approach to the training and evaluation methods currently in use.
In conclusion, our paper contributes significantly to the field in the following ways:
1. Introduction of a novel evaluation principle and the accompanying open-source benchmark DiagGSM8K.  [^1] We have open-sourced our benchmark and evaluation results in https://github.com/dvlab-research/DiagGSM8K This benchmark challenges language models to engage in meta-reasoning, adding a new dimension to the evaluation of cognitive abilities in LLMs, beyond traditional benchmarks.
2. Comprehensive analysis of several SOTA models using the DiagGSM8K benchmark, highlighting critical shortcomings in the current training and evaluation paradigms. This analysis provides valuable insights into the true cognitive capabilities of these models.
3. Advocacy for a paradigm shift in the design and implementation of benchmarks. Our research emphasizes the need for benchmarks that go beyond surface-level evaluations and probe deeper into the reasoning capabilities of LLMs, fostering more sophisticated and nuanced AI development.
2 Related Works
Complex reasoning tasks like math problems have long been widely accepted as a great proxy to fathom the cognitive ability in language models (Sharples1989ComputersAT; KoncelKedziorski2016MAWPSAM; Szegedy2020APP; Polu2020GenerativeLM; Miao2020ADC; hendrycks2021measuring; cobbe2021training). It demands the ability to understand the symbols and texts behind the problems, to dissect the problems into logically connected sub-ones, then combine and arrange results into final solutions. It touches on the cognitive abilities to induce patterns out of problems, to recall corresponding formulae, apply the rules deductively and reason in abstract symbolic way.
GSM8K (cobbe2021training) and MATH (hendrycks2021measuring) have been two popular benchmarks to evaluate the math reasoning capabilities of LLMs in the past few years. Wei2022ChainOT proposed chain of thought to approach the multi-step reasoning tasks in a step by step manner. cobbe2021training and Uesato2022SolvingMW suggested to train a outcome based verifier model and apply reinforcement learning to help improve the performance of language models on GSM8k. Along the reinforcement learning route, Lightman2023LetsVS and luo2023wizardmath applied process based verifier model to aid in the training and inference stage. Stanford Alpaca (alpaca) popularized the knowledge distillation method of cloning corresponding abilities from ChatGPT (OpenAI2022CHATGPT) by asking it to generate related QA pairs. WizardMath (luo2023wizardmath) enhanced the distillation by specifying the QA difficulties in the generation process. Mammoth (yue2023mammoth) combined chain of thought and program of thought, and finetunes its models with the answer generated by GPT-4 (OpenAI2023GPT4TR) that are either in natural language or code language. MetaMath (yu2023metamath) expanded the generated question types by introducing the forward/backward reasoning variations.
Despite the remarkable progress made in math reasoning and other natural language understanding tasks such as MMLU (Hendrycks2020MeasuringMM) and BBH (Suzgun2022ChallengingBT), there are some evidence that shows large language models might not have really mastered the reasoning and even not understanding what it generated. For example, Dziri2023FaithAF evaluated a few compositional reasoning problems and found that LLMs fail to generalize to questions with different complexity than the questions the model was trained on. West2023TheGA found the paradox that AI models can confidently generate some output that surpass average human level but failed to present basic levels of understanding on the output generated. Arkoudas2023GPT4CR used a collection of 21 reasoning tasks to systematically demonstrated that, despite occasional flashes of analytical brilliance, GPT4 is still severely limited when it comes to reasoning. Huang2023LargeLM and Yen2023ThreeQC also found that ChatGPT is very limited in judging the solution correctness of math problems, however our work focuses on the construction of a qualitative and quantitative evaluation framework and focus on the discussion of the evaluation principle and deficiencies of current training paradigm.
Figure 3: This figure illustrates the structure of the DiagGSM8K benchmark, which incorporates math problems from GSM8K cobbe2021training and introduces two additional variations: Program of Thought (POT) (yue2023mammoth) and backward reasoning (yu2023metamath). These variations serve as sophisticated proxy challenges to assess whether language models possess the ability to ’reason about reasoning.’ Each question-solution pair within the benchmark demands that the evaluated model not only determines the solution’s correctness but also identifies the first erroneous step and articulates the underlying error rationale. Note despite the simplicity of the problem, GPT4 fail to predict the correctness of the POT variation.
3 Evaluation Framework
As introduced in Section1, our hypothesis posits that language models, when trained with the aim of optimizing a limited set of reasoning paths within a vast solution space, may lack the capability to comprehend and diagnose alternative reasoning approaches. To test this hypothesis, we developed the DiagGSM8K benchmark, specifically designed to uncover deficiencies in the current training paradigms. As illustrated in Figure-3, given a GSM8K question and its solution, the evaluated model is tasked to predict the correctness of the solution. If the solution is incorrect, the model is expected to further locate the first error location and elucidate the error reason.
Note that each test problem is combined with two variations which requires code solution (yue2023mammoth) and backward reasoning (yu2023metamath). The variations types are chosen specifically due to their significance in expanding the reasoning methodologies in LLMs. The ’Program of Thought,’ a concept proposed by Madaan2022LanguageMO and empirically proven effective in math reasoning by yue2023mammoth; Gou2023ToRAAT, represents a robust reasoning framework. The reasoning process in code language is inherently more structured and hierarchically abstracted. Its reasoning process is also less prone to error types such as calculation error. Backward reasoning, on the other hand, is recently brought under spotlight with discussions on if language models are able to learn backward relations effectively (Berglund2023TheRC).
For each problem collected, we utilized MetaMath-7B (yu2023metamath) with a temperature setting of 1 to generate step by step solutions. A panel of selected annotators were then recruited to review each QA pair on its reasoning process and decide the solution correctness, first error step and error reason. Followings are the annotation procedures adopted by the annotators:
Solution Correctness: Solutions that yield a final output differing from the established ground truth are automatically marked as incorrect. However, in cases where the solution’s final output aligns with the ground truth, annotators are tasked with reviewing the entire reasoning path. Their objective is to ascertain whether the correct output is the result of a logical and sensible reasoning process.
First Error Step: This attribute is applicable for solutions with either an unmatched final output or a matched final output underpinned by flawed reasoning. Annotators identify the initial step in the reasoning process where the logic deviates from correctness. In line with the approach of Lightman2023LetsVS, we dissected GSM8K solutions into discrete steps, each marked by a newline character, and indexed them accordingly. Each step is then categorized as positive, neutral, or negative. Positive and neutral steps represent stages in the reasoning process where the correct final output remains attainable, whereas negative steps indicate a divergence from the path leading to the correct solution.
Error Analysis: Beyond identifying the first erroneous step, annotators are also responsible for conducting an in-depth analysis of the reasoning that led to the error. This involves an examination of the solution’s reasoning flow, focusing on the cause of the initial error and what the correct line of reasoning should have been at that juncture. This error analysis is subsequently compared against the reasoning errors identified by the evaluated models during testing, to assess their accuracy and validity.
Question Types Correct Incorrect Total
Original 692 726 1418
POT 113 109 222
Reverse 622 738 1360
Total 1427 1573 3000
Table 1: This table presents the composition of the DiagGSM8K benchmark. It categorizes questions into ’Original’ (based on GSM8K), ’POT’ (Program of Thought), and ’Reverse’ types, with counts for both correct and incorrect reasoning processes. Due to the difficulties of labelling coding solutions, the POT types of problems are labelled by the author manually therefore smaller in size.
Table 1 shows the statistics about the evaluation benchmark we curated. For more details of the annotation process such as the process design, annotation format please refer to Appendix-LABEL:sec:appendix_annotation.
4 Experiments
For each question-answer pair annotated, the evaluated model are supposed to decide the correctness of the solution and report the error step and error reason if any. The solution correctness and first error step is scored automatically based on the manual annotation result. Only when the evaluated model correctly identified the incorrect solution and error step will its error reason be further examined manually.
4.1 Commercial LLMs Evaluation
Among the closed-source commercial LLMs, ChatGPT and Claude series are the most well-known ones that have been through comprehensive instruction tuning (ouyang2022training; ANTHROPIC2023CLAUDE) and careful alignment process (OpenAI2023GPT4TR; Bai2022TrainingAH; Ganguli2022RedTL). For this study, we specifically evaluated GPT3-5-turbo-0613, Claude2.0, GPT4-0613. Given that these models have demonstrated over 80 percent accuracy in single-pass settings (luo2023wizardmath), our interest lies in assessing their performance under zero-shot conditions in the DiagGSM8K benchmark.
In line with findings from Orca-2 (Mitra2023Orca2T) and other researchers, we recognized that task performance is significantly influenced by system-instruction design. Thus, we conducted empirical experiments with various instruction templates, selecting the most effective one (see Figure-LABEL:fig:PromptsApis) from our validation set for use across all models. The key insight here is that even though Chain-of-thought helps improving the diagnostic reasoning abilities, the expert tailored diagnosing procedure plays a crucial role. For instance, when models were prompted to solve the questions before contrasting their solutions with the provided answers, we observed a bias towards the provided solutions, adversely affecting diagnostic accuracy. Another key aspect of our experimental setup involves the sampling temperature. Unlike the greedy sampling method used in the aforementioned single-pass experiments, we found that greedy decoding substantially diminishes generation quality and diagnostic performance in DiagGSM8K. Consequently, after several iterations of testing on our validation set, we set the sampling temperature to 0.5 for all models.
Table-2 presents the evaluation results for the three selected models. It’s evident that GPT3-5 trails behind both Claude2 and GPT4 across all metrics. For instance, in binary correctness prediction, GPT3-5 achieves a score of only 56.7%, while Claude2 and GPT4 score 65.6% and 78.6%, respectively. Figure-4 offers a visual representation of the incorrect solutions within DiagGSM8K, elucidating the progressively complex nature of the tasks – identifying incorrect solutions, locating the first error step, and articulating the error’s rationale. The graph reveals a downward trend in accuracy for these successive tasks. Notably, GPT3-5 shows a mere 4.3% success rate in identifying the first error step and explaining the reason behind it.
Further scrutiny of Type I (False Positive Rate, FPR) and Type II (False Negative Rate, FNR) errors made by the models, as illustrated in Figure-5, uncovers distinct tendencies in each model’s performance. Claude2 exhibits a relatively balanced distribution between Type I and Type II errors. In contrast, GPT4 is prone to committing Type II errors at a rate approximately 2.5 times higher than Type I errors, indicating a tendency to incorrectly label correct solutions as incorrect. Conversely, GPT3-5 displays an opposite trend to GPT4, with a high false positive rate of 63.4%. This suggests that GPT3-5 is more inclined to uncritically accept the given solutions, irrespective of their actual correctness, as compared to the other two models. Possible interpretation could be made by observing the low accuracy of GPT3-5 in determining the error reasons. The overall uncertainty towards a solution combined with the insufficient calibration of RLHF potentially lead to the sycophancy behavior (Sharma2023TowardsUS).
Despite the fact that all three evaluated models exhibit relatively similar pass@1 rates on the GSM8K dataset (e.g., GPT4 surpassing Claude2 by only 4 percent in accuracy as per luo2023wizardmath), the performance disparity becomes stark in the context of our DiagGSM8K benchmark. GPT4 demonstrates a significantly higher proficiency, being approximately 10 times more effective than GPT3-5 and 4 times more so than Claude2 in accurately diagnosing incorrect solutions from the same dataset. This pronounced variance in evaluation metrics highlights the limitations inherent in the current design of benchmark datasets, reinforcing the argument presented in Section-1 for the necessity of a new paradigm to comprehensively evaluate the reasoning abilities of contemporary LLMs.
While GPT4 shows a dominant performance in comparison to the other two models, it’s important to note instances where either GPT3-5 or Claude2 successfully identify errors or provide plausible explanations where GPT4 does not. This observation suggests that even the most advanced models have limitations and specific scenarios where their performance can be outmatched. For more examples of the responses from evaluated models, please refer to Appendix-LABEL:sec:appendix_examples.
Model Eval Method Accuracy TPR TNR Step Step+Reason
Claude2 0-shot 1968/3000 962/1427 1006/1573 311/1573 173/1573
GPT3-5 0-shot 1701/3000 1125/1427 576/1573 159/1573 68/1573
GPT4 0-shot 2359/3000 985/1427 1374/1573 784/1573 644/1573
WizardMath-70B 3-shot 1187/3000 1176/1427 11/1573 4/1573 1/1573
Mammoth-70B 3-shot 1451/3000 1410/1427 41/1573 4/1573 1/1573
MetaMath-70B 3-shot 1471/3000 1305/1427 166/1573 22/1573 6/1573
llama2-70B-diag 0-shot 1609/3000 453/1427 1156/1573 323/1573 99/1573
Table 2: Evaluation Results of Models on DiagGSM8K: This table presents a detailed breakdown of each model’s performance, including True Positive Rate (TPR) and True Negative Rate (TNR). The ’Step’ column represents the number of instances where models accurately identified an incorrect solution and pinpointed the first error step. ’Step+Reason’ indicates instances where models not only located the first error step in incorrect solutions but also correctly explained the error’s rationale.
Figure 4: Performance Visualization on Incorrect DiagGSM8K Solutions: This graph depicts the performance of Claude2, GPT3-5, and GPT4 in identifying incorrect solutions in DiagGSM8K. ’TNR’ denotes the True Negative Rate. ’TN+Step’ refers to the ratio where models correctly identified an incorrect solution and located the first error step. ’TN+Step+Reason’ represents the frequency of models correctly determining solution correctness, identifying the first error step, and providing an accurate error reason.
Figure 5: Error Rate Analysis of Closed-Source Models on DiagGSM8K: This figure illustrates the True Positive Rate (TPR), True Negative Rate (TNR), False Positive Rate (FPR), and False Negative Rate (FNR) of closed-source models evaluated on DiagGSM8K, offering insights into their accuracy in various aspects of solution evaluation.
4.2 Open-sourced LLMs Evaluation
As mentioned briefly in Section 1 and 2, recent works of fine-tuning language models with program of thought, different reasoning types, and different difficulties have all proven effective in enhancing GSM8K benchmark results. However, just like GPT3-5 has similar pass@1 rate with GPT4 in GSM8K test set, will these open-sourced models specialized in math reasoning performs differently in the DiagGSM8K?
To address this query, we selected several state-of-the-art open-source models fine-tuned on the 70B Llama architecture (touvron2023llama). Namely, they are WizardMath-70B(luo2023wizardmath), Mammonth-70B(yue2023mammoth) and MetaMath-70B(yu2023metamath). WizardMath-70B underwent training on augmented problems of varying difficulty levels and was further refined through a process-oriented reinforcement learning procedure. Mammoth-70B received training on a diverse range of STEM-related problems, including those from the GSM8K dataset, which were augmented with GPT4-generated code solutions. MetaMath-70B’s training encompassed a substantial volume of augmented data, including rephrased and backward-transformed problems from the GSM8K set. Given that these models are specialized and not fine-tuned for general instructions, we employed three-shot in-context learning examples during inference to guide the models in adhering to the desired format and reasoning logic.
As demonstrated in Table-2, despite having exposure to similar training data, all three models failed miserably in the DiagGSM8K benchmark. This outcome highlights their lack of generalization capabilities when faced with problems similar to their training data but presented in different formats. Among these models, MetaMath displayed the most commendable performance, which is not entirely unexpected. Over half of its training dataset (approximately 240k instances) is derived directly from the GSM8K dataset, including a balanced mix of answer augmentations, question rephrasings, and backward transformations of questions. Despite having a training dataset with question types akin to those in DiagGSM8K (e.g., original and POT questions), the Mammoth model failed to demonstrate a deeper meta-understanding of the data, correctly answering only one out of 1573 incorrect solutions. WizardMath, trained using the Proximal Policy Optimization algorithm (Schulman2017ProximalPO) to optimize the joint reward of instruction adherence and solution process, also underperformed on our benchmark.
4.3 In Domain Finetuning
Given the challenges posed by the novel ’reason about reasoning’ task paradigm, a pertinent question arises: Can targeted task-specific training data enhance the performance of current state-of-the-art (SOTA) models on this task? To investigate this, we considered augmenting the GSM8K training set with diagnostics in a similar format. However, due to the labor-intensive nature of manually annotation we opted for a more feasible approach. Using an expert-designed procedure (as demonstrated in Figure-LABEL:fig:TrainingGeneration), we employed GPT-4 to generate the training data. This process involved presenting GPT-4 with a question and its correct solution, then instructing it to introduce an error at a randomly chosen step and complete the solution accordingly. The step-by-step analysis was subsequently generated, focusing on the fabricated error. Despite GPT-4’s modest 40% accuracy in correctly identifying incorrect solutions in the test set, this procedure successfully generated accurate diagnostic training data with over 90 percent accuracy, as verified by expert annotation on a sub-sampled set. It’s important to note that this training data was based solely on the original GSM8K problems, without incorporating any Program of Thought (POT) or backward transformations.
For our base model, we utilized Llama-2-70B-base, aligning with the approach of other open-source SOTA math models. We merged the GSM8K training set with the GPT-4 generated diagnostic data that composed of 5k incorrect solutions and 4k correct solutions. For fine-tuning, we adopted the Qlora method (Dettmers2023QLoRAEF), maintaining the same hyperparameters as used for MetaMath-70B. The outcomes of this approach are detailed in the last row of Table-2. Remarkably, despite the zero-shot setting and a considerably smaller training set, the fine-tuned Llama2 model outperformed all open-source models and even surpassed GPT3-5. Notably, the fine-tuned Llama2 model demonstrated a tendency distinct from that of GPT3-5 and other open-source models; it was less inclined to accept solutions but to over reject solutions regardless of the correctness. As depicted in Figure-6, of the 99 questions where the model accurately predicted both correctness and the first error step, a significant portion involved questions with POT and backward reasoning types. This is particularly noteworthy given that the model was trained exclusively on original questions.
Figure 6: Problem types of incorrect solutions that llama2-70B-diag models has successfully found error step and error reason. Note the training set only includes solutions from original problem.
Caution is necessary when interpreting the outcomes of in-domain fine-tuning. Although the fine-tuned model achieved results comparable to GPT3-5 post-fine-tuning, it’s important to note that the overall number of correct diagnoses for incorrect solutions remains relatively low. This underscores the challenging nature of our DiagGSM8K benchmark, where effective diagnosis across diverse solution spaces requires a comprehensive understanding of the problem. Consequently, simple fine-tuning strategies may not yield substantial improvements in performance. Another significant observation, as indicated in Table-3, is that despite outperforming GPT3-5 in DiagGSM8K, the model’s accuracy on the GSM8K test-set still trails behind that of GPT3-5. This discrepancy could be partially attributed to a lower True Positive Rate in correct solutions and a reduced efficacy in articulating error reasons for incorrect solutions. The fine-tuning process may have enabled the model to replicate diagnostic behaviors, but this approach does not seem to enhance the model’s fundamental comprehension of mathematical reasoning. Another interpretation worth considering is that language models, to a large extent, continue to function as ’stochastic parrots’ (Bender2021OnTD), lacking genuine understanding of the content they generate. It would be intriguing to explore how scaling up the diagnostic data or employing more sophisticated training methodologies might alter these outcomes. We will delve further into this discussion in the following section.
Models GSM8k Diag-SR
GPT4 92.0% 644/1573
Claude2 88.0% 173/1573
GPT3-5 80.8% 68/1573
WizardMath-70B 81.6% 1/1573
Mammoth-70B 76.7% 1/1573
MetaMath-70B 82.3% 6/1573
llama2-70B-GSM8k 74.9% N/A
llama2-70B-diag 74.3% 99/1573
Table 3: Comparison of the performances of SOTA models on GSM8K and DiagGSM8K. The GSM8K results are retrieved from corresponding paper. Diag-SR stands for the correct ratio that models successfully found the erorr step and error reason for the incorrect solutions. llama2-70B-GSM8K are llama2-70B model finetuned on the training set of GSM8k only. llama2-70B-diag are llama2-70B model finetuned on the GSM8k training set and its diagnosis augmentation by GPT4.
5 Discussion and Limitations
Exploring the ’Reason About Reasoning’ Paradigm: A Robust Indicator of Reasoning Capability and Conceptual Mastery?
In fine-tuning models for any given task, the goal often extends beyond achieving a good fit within the training distribution; it also encompasses attaining substantial generalization to unseen instances. This is particularly crucial in complex reasoning tasks, where the focus transcends mere performance metrics on a test set. What becomes equally important is the validity and logic of the reasoning process employed by the evaluated model. Consider this analogy: Understanding the radix/base system in numerals should make operations like addition in base-10 or base-9 similarly challenging. Likewise, for individuals adept at using scratchpads for calculations, the complexity of multiplying two-digit versus three-digit numbers shouldn’t differ drastically. Yet, instances of GPT-4 failing to generalize in such scenarios (Wu2023ReasoningOR; Dziri2023FaithAF) point towards a deficiency in the underlying mastery of these concepts, a flaw that mere accuracy metrics in conventional benchmarks may fail to reveal. This is exacerbated by the current LLMs’ ability to produce well-structured chain of thought responses, giving an illusion of proficiency in various reasoning tasks despite potential gaps in understanding.
This observation underscores a key issue: an intelligent agent that aligns epistemically with human reasoning is expected to base its logic on abstracted rules and principles and apply these deductively across varied contexts. Therefore, the ’reason about reasoning’ paradigm emerges as a vital meta-evaluative tool. It allows us to delve deeper into the reasoning process itself, probing the model’s ability to apply learned concepts flexibly and logically across different scenarios, and unmasking superficial understandings that might be concealed by high performance in traditional test settings.
Limitations of the ’Reason About Reasoning’ Evaluation Paradigm and DiagGSM8K Dataset
Reflecting on Goodhart’s law, which states that ’When a measure becomes a target, it ceases to be a good measure,’ it’s evident that the ’reason about reasoning’ paradigm is not immune to this phenomenon. This paradigm, like any other, can be targeted for optimization, as illustrated in Section-4.3. This is particularly pertinent for static benchmarks where in-domain augmentation and overfitting are feasible. Nonetheless, we contend that our evaluation paradigm presents a greater challenge to overfitting compared to others, owing to its demand for a comprehensive understanding of the problem within a broad error space. While the current DiagGSM8K benchmark is constructed from incorrect solutions generated by the MetaMath-7B model, its scope for expansion is vast. Future versions could incorporate solutions from humans, various models, and even across different languages, enriching its complexity and utility.
Utilizing DiagGSM8K for Model Evaluation Without Manual Labelling
The complexity of the diagnostic process in DiagGSM8K necessitated manual labelling of error reasons in this study to ensure accuracy. However, it is conceivable to use GPT4 for scoring error reasons, with the evaluation of solution correctness and first error step being automated. Table-LABEL:tab:confusionMatrix displays a confusion matrix based on GPT4’s labelling of 100 sampled evaluation instances. While GPT4 exhibits a high true-positive rate, it also has a notable false-positive rate. Its overall accuracy in diagnosing incorrect error reasons aligns closely with its performance in providing error reasons, as shown in Figure-4. Consequently, we advise caution when using GPT4 for automated evaluation. Furthermore, we encourage large tech companies, which have the resources to bear manual labelling costs, to release open-source manual labelling results when publishing findings using DiagGSM8k.