### 20 Jan 2024 v2

# MR-GSM8K: A Meta-Reasoning Revolution in Large Language Model Evaluation

Abstract
In this work, we introduce a novel evaluation paradigm for Large Language Models, one that challenges them to engage in meta-reasoning. This approach addresses critical shortcomings in existing math problem-solving benchmarks, traditionally used to evaluate the cognitive capabilities of agents. Our paradigm shifts the focus from result-oriented assessments, which often overlook the reasoning process, to a more holistic evaluation that effectively differentiates the cognitive capabilities among models. For example, in our benchmark, GPT-4 demonstrates a performance five times better than GPT3-5. The significance of this new paradigm lies in its ability to reveal potential cognitive deficiencies in LLMs that current benchmarks, such as GSM8K, fail to uncover due to their saturation and lack of effective differentiation among varying reasoning abilities. Our comprehensive analysis includes several state-of-the-art math models from both open-source and closed-source communities, uncovering fundamental deficiencies in their training and evaluation approaches. This paper not only advocates for a paradigm shift in the assessment of LLMs but also contributes to the ongoing discourse on the trajectory towards Artificial General Intelligence (AGI). By promoting the adoption of meta-reasoning evaluation methods similar to ours, we aim to facilitate a more accurate assessment of the true cognitive abilities of LLMs.
MR-GSM8K: A Meta-Reasoning Revolution in Large Language Model Evaluation
Zhongshen Zeng Chinese University of Hong Kong zszeng23@cse.cuhk.edu.hk                        Pengguang Chen Chinese University of Hong Kong pgchen@cse.cuhk.edu.hk
Shu Liu Smartmore Co.Ltd sliu@smartmore.com                        Haiyun Jiang thanks: Corresponding author. Tencent AI Lab haiyunjiang@tencent.com                        Jiaya Jia thanks: Corresponding author. Chinese University of Hong Kong leojia@cse.cuhk.edu.hk
1 Introduction
Pretrained on trillions of tokens and possessed with billions of parameters, today’s large language model (OpenAI, 2023; Anthropic, 2023; Touvron et al., 2023) is capable of generating coherent texts that are grammatically correct and seemingly meaningful. If further enhanced with the instruction tuning and human values alignment (e.g. through reinforcement learning from human feedback Ouyang et al., 2022), their outputs are on par with and even more superior than that of human beings. With the hope of differentiating different model’s cognitive ability, math questions are often selected as a proxy evaluation task. However, despite the complexity and diversity of these math problems, recent SOTA LLMs (OpenAI, 2023; Yu et al., 2023; Gou et al., 2023) have been able to achieve accuracy rates exceeding 80% (Luo et al., 2023) on multi-step math reasoning datasets like GSM8K (Cobbe et al., 2021).
Upon a detailed examination of the design principles and objectives of current math datasets, we identified several key shortcomings. Firstly, the majority of these datasets focus on result-oriented metrics, scoring accuracy based solely on the final answer, without considering the underlying reasoning process. With the chain of thought methodology (Wei et al., 2022) and its derivative techniques (Chen et al., 2022; Yao et al., 2023) emerged as the de facto standard for reasoning processes, we argue that the result-driven evaluation method may be insufficient for a comprehensive assessment of the intended cognitive and reasoning capabilities. Secondly, as a recent study (Paster, 2023) suggests that some LLMs who achieved SOTA performances in GSM8K and MATH (Hendrycks et al., 2021) benchmarks have unexpectedly low performance when facing newly released Hungarian high school exams. This raises concerns about the data contamination and potential overfitting to the benchmarks. It also challenges the efficacy of these benchmarks in differentiating model capabilities.
Figure 1: Comparative Performance in GSM8K and MR-GSM8k: This graph illustrates the performance of GPT-4, Claude2, and GPT3-5 in the standard GSM8K benchmark versus our novel MR-GSM8k benchmark. While these models show similar performance levels in GSM8K, a marked variance is observed in MR-GSM8k, highlighting the differentiating power of our benchmark in evaluating deeper reasoning capabilities.
Figure 2: Analysis of Training Sets and Reasoning Path Mastery: The left side of the figure demonstrates the typical inclusion of correct reasoning paths in current math reasoning training sets, despite variations focusing on aspects such as reasoning types (e.g. program of thought) or solution diversity. The right side depicts the limitations of models trained exclusively on correct reasoning paths, showing their inability to assess the validity of alternative reasoning paths for the same problem. This highlights a critical gap in the training paradigm, where models demonstrate basic imitation skills but lack a deep understanding of the underlying logical rules, leading to a superficial grasp of reasoning processes.
In response to these identified limitations, we have introduced a novel paradigm, namely challenging LLMs to reason about different reasoning. This approach is specifically designed to augment existing evaluation benchmarks (as illustrated in Figure-3). Under this paradigm, LLMs are tasked to adopt a role akin to that of a teacher, assessing solutions by determining correctness, identifying potential initial errors, and providing reasons for these errors. Following this design principle, we have developed a novel benchmark named Meta-Reasoning-GSM8k (e.g.MR-GSM8k). This benchmark, characterized by instances manually labeled by experts and rigorously reviewed, serves as a robust tool for a qualitative and quantitative assessment of language models. Our findings indicate that most SOTA models demonstrate a significant performance decline in this more nuanced assessment. As demonstrated in Figure-1, although SOTA models exhibit comparable performance in GSM8K, there is a considerable variance in their effectiveness on our benchmark, with discrepancies up to fivefold.
We argue that our evaluation paradigm does not only provide a more differentiating metric that focuses on the reasoning process over mere computational outcomes, but also sheds light on fundamental deficiencies within the current evaluation and training approaches. As visualized by Figure-2, we observe that recent SOTA models—regardless of whether they are trained on datasets that augment the respective benchmarks in terms of difficulty (Luo et al., 2023), coverage (Lee et al., 2023; Yue et al., 2023), reasoning types (Yu et al., 2023; Yue et al., 2023), solution diversities (Yuan et al., 2023), solution types (Gou et al., 2023; Yue et al., 2023), or through reinforcement learning (Uesato et al., 2022; Luo et al., 2023), are essentially trained by teaching the language models to output a few solution paths within a large search space. Although these models can generate seemingly correct solutions, their grasp of underlying rationale and principles is often superficial and unsophisticated. This is evidenced by issues such as unit inconsistency during calculations (Toh et al., 2023) which highlights a lack of fundamental ontological understanding, and the inability to discern nuanced differences between various reasoning paths for the same problems it was trained on. These core cognitive understanding abilities, vital for benchmarks to assess, have been overlooked due to the prevalent result-oriented metrics and evaluation paradigms. For downstream applications, particularly in education and consulting, the critical factor for success lies in the capability to rigorously analyze different plausible solutions, offering a comprehensive and well-rounded exploration of the solution space. We will provide a more in-depth discussion in Section-5.
The ’reason about reasoning’ principle, central to this benchmark, represents a versatile evaluation approach that can be readily integrated into various existing benchmark suites to enhance their comprehensiveness. Our investigations, detailed in Section-4, uncover significant limitations in the performance of current state-of-the-art models. Furthermore, as discussed in Section-5, we observe that prevalent training paradigms predominantly focus on inductive learning, often failing to cultivate the emergence of underlying rules and principles. This observation underscores the need for a more reflective approach to the training and evaluation methods currently in use.
In conclusion, our paper contributes significantly to the field in the following ways:
1. Introduction of a novel evaluation principle and the accompanying open-source benchmark MR-GSM8k. [^1] We have open-sourced our benchmark and evaluation results in https://github.com/dvlab-research/MR-GSM8KThis benchmark challenges language models to engage in meta-reasoning, adding a new dimension to the evaluation of cognitive abilities in LLMs, beyond traditional benchmarks.
2. Comprehensive analysis of several SOTA models using the MR-GSM8k benchmark, highlighting critical shortcomings in the current training and evaluation paradigms. This analysis provides valuable insights into the true cognitive capabilities of these models.
3. Advocacy for a paradigm shift in the design and implementation of benchmarks. Our research emphasizes the need for benchmarks that go beyond surface-level evaluations and probe deeper into the reasoning capabilities of LLMs, fostering more sophisticated and nuanced AI development.
2 Related Works
Complex reasoning tasks like math problems have long been widely accepted as a great proxy to fathom the cognitive ability in language models (Sharples et al., 1989; Koncel-Kedziorski et al., 2016; Szegedy, 2020; Polu and Sutskever, 2020; Miao et al., 2020; Hendrycks et al., 2021; Cobbe et al., 2021). It demands the ability to understand the symbols and texts behind the problems, to dissect the problems into logically connected sub-ones, then combine and arrange results into final solutions. It touches on the cognitive abilities to induce patterns out of problems, to recall corresponding formulae, apply the rules deductively and reason in abstract symbolic way.
GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) have been two popular benchmarks to evaluate the math reasoning capabilities of LLMs in the past few years. Wei et al., 2022 proposed chain of thought to approach the multi-step reasoning tasks in a step by step manner. Cobbe et al., 2021 and Uesato et al., 2022 suggested to train a outcome based verifier model and apply reinforcement learning to help improve the performance of language models on GSM8k. Along the reinforcement learning route, Lightman et al., 2023 and Luo et al., 2023 applied process based verifier model to aid in the training and inference stage. Stanford Alpaca (Taori et al., 2023) popularized the knowledge distillation method of cloning corresponding abilities from ChatGPT (OpenAI, 2022) by asking it to generate related QA pairs. WizardMath (Luo et al., 2023) enhanced the distillation by specifying the QA difficulties in the generation process. Mammoth (Yue et al., 2023) combined chain of thought and program of thought, and finetunes its models with the answer generated by GPT-4 (OpenAI, 2023) that are either in natural language or code language. MetaMath (Yu et al., 2023) expanded the generated question types by introducing the forward/backward reasoning variations.
Despite the remarkable progress made in math reasoning and other natural language understanding tasks such as MMLU (Hendrycks et al., 2020) and BBH (Suzgun et al., 2022), there are some evidence that shows large language models might not have really mastered the reasoning and even not understanding what it generated. For example, Dziri et al., 2023 evaluated a few compositional reasoning problems and found that LLMs fail to generalize to questions with different complexity than the questions the model was trained on. West et al., 2023 found the paradox that AI models can confidently generate some output that surpass average human level but failed to present basic levels of understanding on the output generated. Arkoudas, 2023 used a collection of 21 reasoning tasks to systematically demonstrated that, despite occasional flashes of analytical brilliance, GPT4 is still severely limited when it comes to reasoning. Huang et al., 2023a and Yen and Hsu, 2023 also found that ChatGPT is very limited in judging the solution correctness of math problems, however our work focuses on the construction of a qualitative and quantitative evaluation framework and focus on the discussion of the evaluation principle and deficiencies of current training paradigm.
Figure 3: This figure illustrates the structure of the MR-GSM8k benchmark, which incorporates math problems from GSM8K Cobbe et al. (2021) and introduces two additional variations: Program of Thought (POT) (Yue et al., 2023) and backward reasoning (Yu et al., 2023). These variations serve as sophisticated proxy challenges to assess whether language models possess the ability to ’reason about reasoning.’ Each question-solution pair within the benchmark demands that the evaluated model not only determines the solution’s correctness but also identifies the first erroneous step and articulates the underlying error rationale. Note despite the simplicity of the problem, GPT4 fail to predict the correctness of the POT variation.
3 Evaluation Framework
3.1 Dataset Construction
As introduced in Section1, our hypothesis posits that language models, when trained with the aim of optimizing a limited set of reasoning paths within a vast solution space, may lack the capability to comprehend and diagnose alternative reasoning approaches. To test this hypothesis, we developed the MR-GSM8k benchmark, specifically designed to uncover deficiencies in the current training paradigms. As illustrated in Figure-3, given a GSM8K question and its solution, the evaluated model is tasked to predict the correctness of the solution. If the solution is incorrect, the model is expected to further locate the first error location and elucidate the error reason.
Note that each test problem is combined with two variations which requires code solution (Yue et al., 2023) and backward reasoning (Yu et al., 2023). The variations types are chosen specifically due to their significance in expanding the reasoning methodologies in LLMs. The ’Program of Thought,’ a concept proposed by Madaan et al., 2022 and empirically proven effective in math reasoning by Yue et al., 2023; Gou et al., 2023, represents a robust reasoning framework. The reasoning process in code language is inherently more structured and hierarchically abstracted. Its reasoning process is also less prone to error types such as calculation error. Backward reasoning, on the other hand, is recently brought under spotlight with discussions on if language models are able to learn backward relations effectively (Berglund et al., 2023).
For each problem collected, we utilized MetaMath-7B (Yu et al., 2023) with a temperature setting of 1 to generate step by step solutions. A panel of selected annotators were then recruited to review each QA pair on its reasoning process and decide the solution correctness, first error step and error reason. Table 1 shows the statistics about the evaluation benchmark we curated. For more details of the annotation process such as the process design, annotation challenges please refer to Appendix-A.
Question Types Correct Incorrect Total
Original 692 726 1418
POT 113 109 222
Reverse 622 738 1360
Total 1427 1573 3000
Table 1: This table presents the composition of the MR-GSM8k benchmark. It categorizes questions into ’Original’ (based on GSM8K), ’POT’ (Program of Thought), and ’Reverse’ types, with counts for both correct and incorrect reasoning processes. Due to the difficulties of labelling coding solutions, the POT types of problems are labelled by the author manually therefore smaller in size.
3.2 Evaluation Metric
For each question-answer pair annotated, the evaluated model are supposed to decide the correctness of the solution and report the first error step and error reason if any. The solution correctness and first error step is scored automatically based on the manual annotation result. Only when the evaluated model correctly identified the incorrect solution and error step will its error reason be further examined manually or automatically by models. Therefore in order to provide a unified and normalized score to reflect the overall competence of the evaluated model, we hereby propose a novel metric named MR-Score.
MR-Score is a weighted combination of three metrics. The first one is the Matthews Correlation Coefficient (e.g. MCC) for the binary classification of solution correctness. The MCC score ranges from -1 to +1 with -1 means total disagreement between prediction and observation, 0 indicates near random performance and +1 represents perfect prediction. Here we interpret negative values as no better than random guess and set 0 as cut-off threshold for normalization purpose. The second metric is the ratio between numbers of solutions with correct first error step predicted and the total number of incorrect solutions. The third metrics is likewise the ratio between number of solutions with correct first error step plus correct error reason predicted and the total number of incorrect solutions. The formula of the MR-Score is then given by
$Score=w_{1}*\max(0,MCC)+w_{2}*ACC_{step}$ (1)
$\quad+w_{3}*ACC_{reason}$
where
$MCC=\frac{TP\times TN-FP\times FN}{\sqrt{(TP+FP)\times(TP+FN)}}$ (2)
×$\quad\times\frac{1}{\sqrt{(TN+FP)\times(TN+FN)}}$
$\begin{split}\text{ACC}_{\text{step}}=\frac{NUM_{\text{correct\_first\_error\_step}}}{NUM_{\text{incorrect\_sols}}}\end{split}$ (3)
$\begin{split}\text{ACC}_{\text{reason}}=\frac{NUM_{\text{correct\_error\_reason}}}{NUM_{\text{incorrect\_sols}}}\end{split}$ (4)
For the weights $w_{1},w_{2}$ and $w_{3}$, they are chosen empirically to be 0.2, 0.3 and 0.5 by considering the difficulties of binary solution correctness classification task, multi-class first error step predictions and free-form error reason explanations. For extended discussion on the design of MR-Score, please check out Appendix-B.
4 Experiments
4.1 Commercial LLMs Evaluation
Among the closed-source commercial LLMs, ChatGPT and Claude series are the most well-known ones that have been through comprehensive instruction tuning (Ouyang et al., 2022; Anthropic, 2023) and careful alignment process (OpenAI, 2023; Bai et al., 2022; Ganguli et al., 2022). For this study, we specifically evaluated GPT3-5-turbo-0613, Claude2.0, GPT4-0613. Given that these models have demonstrated over 80 percent accuracy in single-pass settings (Luo et al., 2023), our interest lies in assessing their performance under zero-shot conditions in the MR-GSM8k benchmark.
In line with findings from Orca-2 (Mitra et al., 2023) and other researchers, we recognized that task performance is significantly influenced by system-instruction design. Thus, we conducted empirical experiments with various instruction templates, selecting the most effective one (see Figure-LABEL:fig:PromptsApis) from our validation set for use across all models. The key insight here is that even though Chain-of-thought helps improving the diagnostic reasoning abilities, the expert tailored diagnosing procedure plays a crucial role. For instance, when models were prompted to solve the questions before contrasting their solutions with the provided answers, we observed a bias towards the provided solutions, adversely affecting diagnostic accuracy. Another key aspect of our experimental setup involves the sampling temperature. Unlike the greedy sampling method used in the aforementioned single-pass experiments, we found that greedy decoding substantially diminishes generation quality and diagnostic performance in MR-GSM8k. Consequently, after several iterations of testing on our validation set, we set the sampling temperature to 0.5 for all models.
Table-2 presents the evaluation results for the three selected models. It’s evident that GPT3-5 trails behind both Claude2 and GPT4 across all metrics. For instance, in binary correctness prediction, GPT3-5 achieves an MCC score of only 0.198, while Claude2 and GPT4 score 0.345 and 0.614, respectively. Figure-4 offers a visual representation of the incorrect solutions within MR-GSM8k, elucidating the progressively complex nature of the tasks – identifying incorrect solutions, locating the first error step, and articulating the error’s rationale. The graph reveals a downward trend in accuracy for these successive tasks. Notably, GPT3-5 shows a mere 4.64% success rate in identifying the first error step and explaining the reason behind it.
Further scrutiny of Type I (False Positive Rate, FPR) and Type II (False Negative Rate, FNR) errors made by the models, as illustrated in Figure-5, uncovers distinct tendencies in each model’s performance. Claude2 exhibits an impressively balanced distribution between Type I and Type II errors. In contrast, GPT4 is prone to committing Type II errors at a rate approximately three times higher than Type I errors, indicating a tendency to incorrectly label correct solutions as incorrect. Conversely, GPT3-5 displays an opposite trend to GPT4, with a high false positive rate of 60.52%. This suggests that GPT3-5 is more inclined to uncritically accept the given solutions, irrespective of their actual correctness, as compared to the other two models. Possible interpretation could be made by observing the low accuracy of GPT3-5 in determining the error reasons. The overall uncertainty towards a solution combined with the potentially insufficient calibration of RLHF lead to the sycophancy behavior (Sharma et al., 2023).
Despite the fact that all three evaluated models exhibit relatively similar pass@1 rates on the GSM8K dataset (e.g., GPT4 surpassing Claude2 by only 4 percent in accuracy as per Luo et al., 2023), the performance disparity becomes stark in the context of our MR-GSM8k benchmark. GPT4 demonstrates a significantly higher proficiency, being approximately 5 times more effective than GPT3-5 and 2.5 times more so than Claude2 measured by MR-Score. This pronounced variance in evaluation metrics highlights the limitations inherent in the current design of benchmark datasets, reinforcing the argument presented in Section-1 for the necessity of a new paradigm to comprehensively evaluate the reasoning abilities of contemporary LLMs.
While GPT4 shows a dominant performance in comparison to the other two models, it’s important to note instances where either GPT3-5 or Claude2 successfully identify errors or provide plausible explanations where GPT4 does not. This observation suggests that even the most advanced models have limitations and specific scenarios where their performance can be outmatched. For more examples of the responses from evaluated models, please refer to Appendix-LABEL:sec:appendix_examples.
Table 2: Evaluation Results of Models on MR-GSM8k: This table presents a detailed breakdown of each model’s performance, including True Positive Rate (TPR), True Negative Rate (TNR) and Matthews Correlation Coefficient. The ’Step’ column represents the accuracy of correctly identifying an incorrect solution and pinpointing the first error step. ’S+R/M’ column showcased the accuracy of not only locating the first error step in incorrect solutions but also correctly explaining the error’s rationale. The overall MR-Score/M is a normalized metric ranges from 0 to 1 and calculated based on formula described in Section-3. The M here means the error reason is labelled manually and MR-Score is calculated based on the results of that.
Model Eval Method TPR TNR MCC Step S+R/M MR-Score/M
Claude2 0-shot 67.41% 67.13% 0.345 21.04% 11.76% 0.191
GPT3-5 0-shot 78.84% 39.48% 0.198 11.38% 4.64% 0.097
GPT4 0-shot 69.03% 90.59% 0.614 52.32% 43.04% 0.495
WizardMath-70B 3-shot 82.41% 2.73% -0.250 0.38% 0.06% 0.001
Mammoth-70B 3-shot 98.81% 2.73% 0.055 0.25% 0.06% 0.012
MetaMath-70B 3-shot 91.45% 10.55% 0.034 1.40% 0.38% 0.013
llama2-70B-diag 0-shot 31.74% 73.49% 0.058 20.79% 6.29% 0.105
Table 2: Evaluation Results of Models on MR-GSM8k: This table presents a detailed breakdown of each model’s performance, including True Positive Rate (TPR), True Negative Rate (TNR) and Matthews Correlation Coefficient. The ’Step’ column represents the accuracy of correctly identifying an incorrect solution and pinpointing the first error step. ’S+R/M’ column showcased the accuracy of not only locating the first error step in incorrect solutions but also correctly explaining the error’s rationale. The overall MR-Score/M is a normalized metric ranges from 0 to 1 and calculated based on formula described in Section-3. The M here means the error reason is labelled manually and MR-Score is calculated based on the results of that.
Figure 4: Performance Visualization on Incorrect MR-GSM8k Solutions: This graph depicts the performance of Claude2, GPT3-5, and GPT4 in identifying incorrect solutions in MR-GSM8k. ’TNR’ denotes the True Negative Rate. ’TN+Step’ refers to the ratio where models correctly identified an incorrect solution and located the first error step. ’TN+Step+Reason’ represents the frequency of models correctly determining solution correctness, identifying the first error step, and providing an accurate error reason.
Figure 5: Error Rate Analysis of Closed-Source Models on MR-GSM8k: This figure illustrates the True Positive Rate (TPR), True Negative Rate (TNR), False Positive Rate (FPR), and False Negative Rate (FNR) of closed-source models evaluated on MR-GSM8k, offering insights into their accuracy in various aspects of solution evaluation.
4.2 Open-sourced LLMs Evaluation
As mentioned briefly in Section 1 and 2, recent works of fine-tuning language models with program of thought, different reasoning types, and different difficulties have all proven effective in enhancing GSM8K benchmark results. However, just like GPT3-5 has similar pass@1 rate with GPT4 in GSM8K test set, will these open-sourced models specialized in math reasoning performs differently in the MR-GSM8k?
To address this query, we selected several state-of-the-art open-source models fine-tuned on the 70B Llama architecture (Touvron et al., 2023). Namely, they are WizardMath-70B(Luo et al., 2023), Mammonth-70B(Yue et al., 2023) and MetaMath-70B(Yu et al., 2023). WizardMath-70B underwent training on augmented problems of varying difficulty levels and was further refined through a process-oriented reinforcement learning procedure. Mammoth-70B received training on a diverse range of STEM-related problems, including those from the GSM8K dataset, which were augmented with GPT4-generated code solutions. MetaMath-70B’s training encompassed a substantial volume of augmented data, including rephrased and backward-transformed problems from the GSM8K set. Given that these models are specialized and not fine-tuned for general instructions, we employed three-shot in-context learning examples during inference to guide the models in adhering to the desired format and reasoning logic.
As demonstrated in Table-2, despite having exposure to similar training data, all three models failed miserably in the MR-GSM8k benchmark. This outcome highlights their lack of generalization capabilities when faced with problems similar to their training data but presented in different formats. Among these models, MetaMath displayed the most commendable performance, which is not entirely unexpected. Over half of its training dataset (approximately 240k instances) is derived directly from the GSM8K dataset, including a balanced mix of answer augmentations, question rephrasings, and backward transformations of questions. Despite having a training dataset with question types akin to those in MR-GSM8k (e.g., original and POT questions), the Mammoth model failed to demonstrate a deeper meta-understanding of the data, correctly answering only one out of 1573 incorrect solutions. WizardMath, trained using the Proximal Policy Optimization algorithm (Schulman et al., 2017) to optimize the joint reward of instruction adherence and solution process, also underperformed on our benchmark.
4.3 In Domain Finetuning
Given the challenges posed by the novel ’reason about reasoning’ task paradigm, a pertinent question arises: Can targeted task-specific training data enhance the performance of current state-of-the-art (SOTA) models on this task? To investigate this, we considered augmenting the GSM8K training set with diagnostics in a similar format. However, due to the labor-intensive nature of manually annotation we opted for a more feasible approach. Using an expert-designed procedure (as demonstrated in Figure-LABEL:fig:TrainingGeneration), we employed GPT-4 to generate the training data. This process involved presenting GPT-4 with a question and its correct solution, then instructing it to introduce an error at a randomly chosen step and complete the solution accordingly. The step-by-step analysis was subsequently generated, focusing on the fabricated error. Despite GPT-4’s modest 40% accuracy in correctly identifying incorrect solutions in the test set, this procedure successfully generated accurate diagnostic training data with over 90 percent accuracy, as verified by expert annotation on a sub-sampled set. It’s important to note that this training data was based solely on the original GSM8K problems, without incorporating any Program of Thought (POT) or backward transformations.
For our base model, we utilized Llama-2-70B-base, aligning with the approach of other open-source SOTA math models. We merged the GSM8K training set with the GPT-4 generated diagnostic data that composed of 5k incorrect solutions and 4k correct solutions. For fine-tuning, we adopted the Qlora method (Dettmers et al., 2023), maintaining the same hyperparameters as used for MetaMath-70B. The outcomes of this approach are detailed in the last row of Table-2. Remarkably, despite the zero-shot setting and a considerably smaller training set, the fine-tuned Llama2 model outperformed all open-source models and even surpassed GPT3-5. Notably, the fine-tuned Llama2 model demonstrated a tendency distinct from that of GPT3-5 and other open-source models; it was less inclined to accept solutions but to over reject solutions regardless of the correctness. As depicted in Figure-6, of the 99 questions where the model accurately predicted both correctness and the first error step, a significant portion involved questions with POT and backward reasoning types. This is particularly noteworthy given that the model was trained exclusively on original questions.
Figure 6: Problem types of incorrect solutions that llama2-70B-MR models has successfully found error step and error reason. Note the training set only includes solutions from original problem.
Caution is necessary when interpreting the outcomes of in-domain fine-tuning. Although the fine-tuned model achieved results comparable to GPT3-5 post-fine-tuning, it’s important to note that the overall number of correct diagnoses for incorrect solutions remains relatively low (e.g. 6.29%). This underscores the challenging nature of our MR-GSM8k benchmark, where effective diagnosis across diverse solution spaces requires a comprehensive understanding of the problem. Consequently, simple fine-tuning strategies may not yield substantial improvements in performance. Another significant observation, as indicated in Table-3, is that despite outperforming GPT3-5 in MR-GSM8k, the model’s accuracy on the GSM8K test-set still trails behind that of GPT3-5. This discrepancy could be partially attributed to a lower True Positive Rate in correct solutions and a reduced efficacy in articulating error reasons for incorrect solutions. The fine-tuning process may have enabled the model to replicate diagnostic behaviors, but this approach does not seem to enhance the model’s fundamental comprehension of mathematical reasoning. Another interpretation worth considering is that language models, to a large extent, continue to function as ’stochastic parrots’ (Bender et al., 2021), lacking genuine understanding of the content they generate. It would be intriguing to explore how scaling up the diagnostic data or employing more sophisticated training methodologies might alter these outcomes. We will delve further into this discussion in the following section.
Models GSM8k MRS/M
GPT4 92.0% 0.495
Claude2 88.0% 0.191
llama2-70B-MR 74.3% 0.105
GPT3-5 80.8% 0.097
MetaMath-70B 82.3% 0.013
Mammoth-70B 76.7% 0.012
WizardMath-70B 81.6% 0.001
llama2-70B-GSM8k 74.9% N/A
Table 3: Comparison of the performances of SOTA models on GSM8K and MR-GSM8k. The GSM8K results are retrieved from corresponding paper. llama2-70B-GSM8K are llama2-70B model finetuned on the training set of GSM8k only. llama2-70B-MR are llama2-70B model finetuned on the GSM8k training set and its meta-reasoning augmentation by GPT4. The MRS/M stand for the MR-Score in MR-GSM8k calculated from manual labelling results.
5 Ethical Discussion and Limitations
Ethical and Societal Dimensions of Meta-Reasoning Evaluation
As briefly discussed in Section-2, the performance on math reasoning tasks are often chosen as a key indicator of the general competence of the evaluated language model, due to its inherent requirements on cognitive abilities such as natural language understanding, concept abstraction, deductive reasoning etc. A common underlying assumption of this logic is that models succeed in reasoning tasks may possess the relevant cognitive/meta capabilities, and such capabilities should remain effective across different domains. Put aside the potential controversies behind this logic, we would like to contend that it is not sufficient to unveil the cognitive depth of the evaluated model by only looking at the computation results. What becomes equally important is the validity and logic of the reasoning process employed by the evaluated model. Therefore, the "reason about reasoning" paradigm emerges as a vital meta-evaluative tool. It allows us to delve deeper into the reasoning process itself, probing the model’s ability to apply learned concepts flexibly and logically across different scenarios.
As illustrated by Figure-2, for the evaluated model to successfully diagnose the solution correctness, it is necessary for the model to be able to infer the correct result and also be able to counterfactually reasons along different reasoning paths and actively examines the conditions and assumptions made on different steps. It is unlikely to succeed in this paradigm without a holistic understanding and robust mastery of the underlying concepts.
Moreover, the meta-reasoning paradigm also opens up a window for the researchers to examine the "alignment" and "grounding" levels of the evaluated model. In subjects of natural science, we would expect an intelligent agent that aligns epistemically with humans to ground its logic on the rules and principles that closely matches the physical world we live in. A quick example would be the unit inconsistency issue during calculations of models (Toh et al., 2023) which highlights a lack of fundamental ontological understanding. In the downstream application scenarios such as education and consultancy, such limitations might present a fundamental deficiency that prevents a larger adoption of LLM technologies.
Limitations of the ’Reason About Reasoning’ Evaluation Paradigm and MR-GSM8k Dataset
Reflecting on Goodhart’s law, which states that ’When a measure becomes a target, it ceases to be a good measure,’ it’s evident that the ’reason about reasoning’ paradigm is not immune to this phenomenon. This paradigm, like any other, can be targeted for optimization, as illustrated in Section-4.3. This is particularly pertinent for static benchmarks where in-domain augmentation and overfitting are feasible. Nonetheless, we contend that our evaluation paradigm presents a greater challenge to overfitting compared to others, owing to its demand for a comprehensive understanding of the problem within a broad error space. While the current MR-GSM8k benchmark is constructed from incorrect solutions generated by the MetaMath-7B model, its scope for expansion is vast. Future versions could incorporate solutions from humans, various models, and even across different languages, enriching its complexity and utility.
Does MR-GSM8k Mandates Human Labelling?
For the sake of rigorousness, every error reason from evaluated models with correct first error step predicted is examined manually in this work. However, this does not imply that MR-GSM8K necessitates a manual labelling for every evaluation. We would like to emphasize that MR-Score is consist of three sub-metrics and error reason is (only) one of the evaluation criteria. Similar to the translation tasks where one expression in one language might corresponds to many variations in another language, it is likewise difficult to come up with an automatic evaluator that scores the error reason perfectly. Nonetheless, this would not undermine the arguments we contend, nor would it affect the cognitive deficiencies unveiled by this metrics. To the best of our knowledge, GPT4 has been the most popular choice for being an automatic evaluator across different metrics (Zheng et al., 2023; Liu et al., 2023). In Appendix-B we empirically demonstrated that GPT4 is able to serve as a decent automatic evaluator that the final MR-Score calculated based on its labelling results is close to that of the manual labelling results.
6 Conclusion
Throughout this paper, we have delved into the inadequacies of prevalent math reasoning benchmarks and introduced a pioneering evaluation paradigm that compels models to engage in meta-reasoning. Our empirical findings showcase that the benchmark, developed under this novel paradigm, stands out in its ability to differentiate between models and uncover their various deficiencies. This has been particularly evident in the struggles of state-of-the-art language models, which, when confronted with our benchmark, have exposed significant shortcomings inherent in the current training methodologies. These revelations advocate a critical reevaluation of existing training and evaluation practices in the realm of large language models.
In advocating for the widespread adoption of our ’reason about reasoning’ evaluation paradigm, we urge researchers to adapt and broaden other reasoning benchmarks in a similar vein. Such transformation is vital not only for a more rigorous assessment of LLMs but also for fostering a deeper and more holistic understanding of these models’ capabilities. By pushing the boundaries of conventional benchmarks and embracing this innovative approach, we can pave the way for more insightful, nuanced evaluations that truly challenge and advance the frontiers of AI research.
Appendix A Annotation Details for MR-GSM8k
In this section we will give a more thorough description of the dataset construction process. We will first present the annotation procedure adopted by our annotators and then describe the challenges faced during such process.
A.1 Annotation Procedure
Given a question and its ground truth solution, the annotator is supposed to first read the question and ground truth solution carefully and make sure he/she comprehend the question thoroughly. Then the annotator is tasked to label the following fields sequentially:
Solution Correctness: Solutions that yield a final output differing from the established ground truth are automatically marked as incorrect. However, in cases where the solution’s final output aligns with the ground truth, annotators are tasked with reviewing the entire reasoning path. Their objective is to ascertain whether the correct output is the result of a logical and sensible reasoning process.
First Error Step: This attribute is applicable for solutions with either an unmatched final output or a matched final output underpinned by flawed reasoning. Annotators identify the initial step in the reasoning process where the logic deviates from correctness. In line with the approach of Lightman et al., 2023, we dissected GSM8K solutions into discrete steps, each marked by a newline character, and indexed them accordingly. Each step is then categorized as positive, neutral, or negative. Positive and neutral steps represent stages in the reasoning process where the correct final output remains attainable, whereas negative steps indicate a divergence from the path leading to the correct solution.
Error Analysis: Beyond identifying the first erroneous step, annotators are also responsible for conducting an in-depth analysis of the reasoning that led to the error. This involves an examination of the solution’s reasoning flow, focusing on the cause of the initial error and what the correct line of reasoning should have been at that juncture. This error analysis is subsequently compared against the reasoning errors identified by the evaluated models during testing, to assess their accuracy and validity.
A.2 Annotation Challenges
The annotation task turns out to be more difficult than we expect at the beginning. The challenge comes from several sources: First, the language barrier hinders the non-native speaker to understand the meaning of the question. Second, the labelling task requires the annotator to read the question, the ground truth solution and the model solution before judging the correctness of the reasoning process. If the reasoning process is problematic, the annotator is further required to find out the first error step and reason about why the model made such mistake. To reason along the path of the model solution and figure out why the error occurred in the perspective of the model and verbalize all the above in the error reason is quite time-consuming and an exhausting task. Third, some questions are ambiguous in its wording and allows for multiple different interpretations. The backward reasoning transformation of the original problem exacerbated this problem.
To help lower the difficulty of labelling, we translated the problems and solutions into Chinese with the help from ChatGPT (OpenAI, 2022). We noticed that ChatGPT occasionally would made some translation errors such as missing critical information and misinterpret the original text, we therefore enclosed both the original problem-solution pairs and the translated ones for reference when in doubt (See Figure-7 for a full example). The solution correctness, first error step and error analysis are then collected from the annotated dictionary and translated back to English.
A.3 Annotation Quality Control
The annotators are selected based on their labelling performance on a balanced small hold-out problems set consist of 50 questions. For every problems in the Mr-GSM8K, they have been through several examinations before used in the evaluation process: First, every question is labelled twice by different annotators. Inconsistent questions will be singled out to be labelled by the quality control labeller. Besides the double annotation, 50 percent of the annotated problems are sampled out for quality control in the second round of verification. During the evaluation process, all the questions that has a matching correctness prediction and error steps are manually examined by the authors of this work for its error reasons. Questions with incorrect error steps or reasons are cleaned up in this final stage.
Figure 7: To lower the requirements and difficulties for labelling, we used ChatGPT to translate the MR-GSM8k problems into Chinese. In case for the potential translation errors (missing information and wrong translation) we included the original English text for reference when in doubt.
Appendix B Design thinking of MR-Score
The MR-Score is consist of three sub-metrics corresponds to the three sequential reasoning sub-tasks. For the first solution correctness prediction, we empirically noticed that most of the evaluated language models tend to blindly classify the given solution as correct, exemplified by the low true-negative rate in Table-2. Therefore, we chose the MCC score instead of metrics like F1 or Balanced-Accuracy due to its value range. The models that have high true-positive rate but low true negative rate will have near zero score under the MCC metric. For the second and third tasks of locating first error step and elucidating error reason, we chose the simple accuracy metric. One of the reason is that locating the first error step is a multi-class classification problem and it is difficult to have large prediction bias while at the same time scores high accuracy. Similarly, the explaining error reason task is a free-form generation task that requires substantial understanding and a simple accuracy metric is enough to categorize the model behavior.
As to the weights given to the three metrics, they are crafted by considering the task complexity and the difference between manual labelling and auto labelling results of the error reason. As discussed in Section-5, we chose the GPT4-Turbo-1106 as our proxy evaluator and Table-4 is the results of auto-labelling vs our expert manual-labelling. It is clear that the final MR-Score calculated from manual labelling VS auto labelling are very close to each other, exhibiting the potential of GPT4 to serve as a delegate evaluator for our task.
Models Step Step+Reason/M Step+Reason/A MR-Score/M MR-Score/A
GPT4 823/1573 677/1573 732/1573 0.495 0.512
Claude2 331/1573 185/1573 224/1573 0.191 0.203
llama2-70B-MR 327/1573 99/1573 139/1573 0.105 0.118
GPT3-5 179/1573 73/1573 73/1573 0.097 0.097
MetaMath-70B 22/1573 6/1573 7/1573 0.013 0.013
Mammoth-70B 4/1573 1/1573 2/1573 0.012 0.012
WizardMath-70B 6/1573 1/1573 1/1573 0.001 0.001
Table 4: Comparison of the manual labelling results and GPT4-Turbo-1106 labelling results. Step column shows the number that each evaluated models successfully located the first error steps among incorrect solutions. Step+Reason/M stands for the manual labelling results of the error reasons where its first error step is correct. Step+Reason/A corresponds to the labelling results of GPT4-Turbo-1106. llama2-70B-MR are llama2-70B model finetuned on the GSM8k training set and its meta-reasoning augmentation by GPT4.
Table-LABEL:tab:confusionMatrix displays the confusion matrix based on GPT4’s labelling of all the error reasons. Notably, GPT4 is able to achieve 82% of overall accuracy despite a substantial higher false positive rate than false negative rate. However, we still encourage large tech companies, who have the resources to bear manual labelling costs, to release open-source manual labelling results when publishing findings using MR-GSM8k for the best of rigorousness.